OpenStudy (anonymous):

if w= -6 find the value of 2w+5

OpenStudy (anonymous):

7

OpenStudy (lilg132):

2(-6) + 5 = -12 + 5 = -7

OpenStudy (anonymous):

-7 *

OpenStudy (lilg132):

remember to click good answer :)

OpenStudy (anonymous):

2(-6)+5= -7

OpenStudy (anonymous):

2(-6)+5= -7